Circular No. 8484

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA CYGNI 2005 H. Naito and N. Tokimasa, Nishi-Harima Astronomical Observatory

(NHAO); and H. Yamaoka, Kyushu University, report: “A low-resolution spectrum (range 400–850 nm, resolution 150) of the possible nova reported on IAUC 8483, obtained on Feb. 11.8 UT with the NHAO 0.60-m telescope, shows prominent and broad Balmer lines in emission, which suggests that it is indeed a classical nova. The FWHM of the Hα line is 3200 km/s. The nearest object in the Digitized Sky Survey images is a star (mag ∼ 18) located 2′′.5 west-southwest of the reported position.”

SUPERNOVA 2005af IN NGC 4945 A. V. Filippenko and R. J. Foley, University of California, Berkeley,

report that inspection of CCD spectra (range 300–920 nm), obtained on Feb. 12 UT with the Keck I 10-m telescope (+ LRIS), shows that the possible supernova in NGC 4945 reported on IAUC 8482 is a supernova of type II, with well-developed P-Cyg profiles of H, Fe II, Ca II, and other species. The general appearance of the spectrum suggests the II-plateau subclass, perhaps 1 month after the explosion.

SUPERNOVA 2005ag Further to IAUC 8478, W. Li reports the LOSS discovery of a super-

nova (mag 18.2) on a KAIT CCD image taken on Feb. 10.55 UT; SN 2005ag is located at α = 14^{h}56^{m}43^{s}.65, δ = +9^{o}19′42′′.5 (equinox 2000.0), which is 6′′.5 east and 5′′.9 north of the center of the apparent host galaxy. Filippenko and Foley add that inspection of noisy CCD spectra, obtained as above, shows that 2005ag is a supernova of type Ia, probably within a week of maximum brightness; the redshift of the host galaxy was not measured.

SUPERNOVAE 2004ht–2004io Further to IAUC 8481, the additional co-authors should be added: J.

Estrada and J. Marriner, Fermilab; A. Becker and G. Miknaitis, University of Washington; and J. Kaplan and R. Romani, Stanford University.

COMET C/2004 Q2 (MACHHOLZ) Corrigendum. On IAUC 8480, lines 8–9, the data for Jan. 15.6344

should read as follows: fan 1, 150′′ in p.a. 282^{o}; fan 2, 150′′ in p.a. 239^{o}; fan 3, 30′′ in p.a. 211^{o}.

2005 February 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green